# Legged Robot State Estimation Using Invariant Neural-Augmented Kalman Filter with a Neural Compensator

### v2

@@ II-A Contact Estimator
where $\mathbf{n}_{i}$ represents the normal vector of the contact surface at contact point $i$.
@@
where $S_{i}$ represents the contact state at contact point $i$.
@@ III-A SE2(3) Group Generation Network
$\xi&=(\omega,v,p)^{\top}\in\mathbb{R}^{9}\\ \sum_{i=1}^{9}\xi_{i}G_{i}&\in\mathfrak{se}(3),\quad\text{where }\xi_{i}=\begin{cases}\omega_{i}&\text{if }i\leq 3,\\ v_{i-3}&\text{if }4\leq i\leq 6,\\ p_{i-6}&\text{if }i>6.\end{cases}\end{aligned}$ (10)
@@
$\text{exp}(\cdot)&:\hat{\eta}_{t}^{r^{\wedge}}\mapsto\hat{\textbf{E}}_{t}^{r}\in\text{SE}_{2}(3)\\ \text{Exp}(\cdot)&:\hat{\eta}_{t}^{r}\mapsto\hat{\textbf{E}}_{t}^{r}\in\text{SE}_{2}(3)\end{aligned}$ (11)
@@ III-B Invariant Neural-Augmented Kalman Filter
where the $\Pi_{t}$ defines the auxiliary selection matrix, which is the $[\textbf{I}\quad\textbf{0}_{3,3}]$, and the gain $\textbf{K}_{t}$ is obtained from (19):
@@
$\textbf{H}_{t}$ is defined to $[\textbf{0}_{3,3}\quad\textbf{0}_{3,3}\quad-\textbf{I}\quad\textbf{I}]$, and $\bar{\textbf{N}}_{t}$ is computed by $\bar{\textbf{R}}_{t}\textbf{J}_{p}(\tilde{q}_{t})\text{Cov}(\textbf{w}_{t}^{q})\textbf{J}_{p}^{\top}(\tilde{q}_{t})\bar{\textbf{R}}_{t}^{\top}$.
@@ IV-A Performance Analysis of SE2(3) Group Generation
The neural network used in this study does not perform operations directly on the Lie group but operates in the Lie algebra, followed by exponential mapping. Since the Lie group is defined on a smooth manifold, performing group operations requires matrix multiplication. In contrast, the Lie algebra, defined in the tangent space, only requires matrix addition for operations. Consequently, obtaining the output in the Lie algebra and applying exponential mapping significantly improves training efficiency. Table I and Fig. 4 show the computation time according to batch size. While matrix addition has a time complexity of O$(n^{2})$, matrix multiplication has a time complexity of O$(n^{3})$. This highlights the advantage of obtaining outputs in the tangent space, where only matrix addition is required, to enhance training efficiency. A difference of approximately 0.1 times in computation time was observed, confirming significantly faster performance.
@@
Figure 5: The visual and numerical results of the trained SEGGN. The first row illustrates the transformed coordinates based on the SEGGN output. The second row presents numerical evaluations, including the Frobenius norm error, calculated as $||\textbf{R}^{\top}\textbf{R}-\textbf{I}||_{F}$, and the determinant of R, both shown as histograms.

### v3

@@ II-A Contact Estimator
where $\mathbf{n}_{i}$ represents the normal vector of the contact surface at contact point ii.
@@
where $S_{i}$ represents the contact state at contact point ii.
@@ III-A SE2(3) Group Generation Network
$\begin{aligned} \xi&=(\omega,v,p)^{\top}\in\mathbb{R}^{9}\\ \sum_{i=1}^{9}\xi_{i}G_{i}&\in\mathfrak{se}(3),\quad\text{where }\xi_{i}=\begin{cases}\omega_{i}&\text{if }i\leq 3,\\ v_{i-3}&\text{if }4\leq i\leq 6,\\ p_{i-6}&\text{if }i>6.\end{cases}\end{aligned}$ (10)
@@
$\begin{aligned} \text{exp}(\cdot)&:\hat{\eta}_{t}^{r^{\wedge}}\mapsto\hat{\textbf{E}}_{t}^{r}\in\text{SE}_{2}(3)\\ \text{Exp}(\cdot)&:\hat{\eta}_{t}^{r}\mapsto\hat{\textbf{E}}_{t}^{r}\in\text{SE}_{2}(3)\end{aligned}$ (11)
@@ III-B Invariant Neural-Augmented Kalman Filter
where the $\Pi_{t}$ defines the auxiliary selection matrix, which is the $\textbf{I}\quad\textbf{0}_{3,3}]$, and the gain $\textbf{K}_{t}$ is obtained from (19):
@@
$\textbf{H}_{t}$ is defined to $\textbf{0}_{3,3}\quad\textbf{0}_{3,3}\quad-\textbf{I}\quad\textbf{I}]$, and $\bar{\textbf{N}}_{t}$ is computed by $\bar{\textbf{R}}_{t}\textbf{J}_{p}(\tilde{q}_{t})\text{Cov}(\textbf{w}_{t}^{q})\textbf{J}_{p}^{\top}(\tilde{q}_{t})\bar{\textbf{R}}_{t}^{\top}$.
@@ IV-A Performance Analysis of SE2(3) Group Generation
[…]
@@
Figure 5: The visual and numerical results of the trained SEGGN. The first row illustrates the transformed coordinates based on the SEGGN output. The second row presents numerical evaluations, including the Frobenius norm error, calculated as $\textbf{R}^{\top}\textbf{R}-\textbf{I}||_{F}$, and the determinant of R, both shown as histograms.